OpenStudy (anonymous):

Given that f(x) = x2 − 3x + 3 and g(x) = the quantity of x minus one, over four, solve for f(g(x)) when x = 5.

OpenStudy (jim766):

substitute the 5 into the inside function....g(x), then take that answer and substitute it into the outside function....f(x)

OpenStudy (anonymous):

is the answer -3 ? @Jim766

OpenStudy (jim766):

f(5) = (x-1)/4 what do you get when you put a 5 in place of the x?

OpenStudy (jim766):

= (5-1)/4

OpenStudy (anonymous):

it = 1? @Jim766

OpenStudy (jim766):

yes....now find f(1)

OpenStudy (jim766):

f(1) = 1^2 - 3(1) + 3

OpenStudy (anonymous):

thats all it asks for. thank you